# A Baseline Method for Removing Invisible Image Watermarks using Deep Image Prior

### v1

@@ Abstract
Image watermarks have been considered a promising technique to help detect AI-generated content, which can be used to protect copyright or prevent fake image abuse. In this work, we present a black-box method for removing invisible image watermarks, without the need of any dataset of watermarked images or any knowledge about the watermark system. Our approach is simple to implement: given a single watermarked image, we regress it by deep image prior (DIP). We show that from the intermediate steps of DIP one can reliably find an evasion image that can remove invisible watermarks while preserving high image quality. Due to its unique working mechanism and practical effectiveness, we advocate including DIP as a baseline invasion method for benchmarking the robustness of watermarking systems. Finally, by showing the limited ability of DIP and other existing black-box methods in evading training-based visible watermarks, we discuss the positive implications on the practical use of training-based visible watermarks to prevent misinformation abuse. Our code is publicly available at https://github.com/HengyueL/DIP_Watermark_Evasion.
@@ 1 Introduction
In this prosperous era of generative AI, the traceability of AI-generated content (e.g., language, images, and videos) to its source has been frequently mentioned as a promising solution to promote the responsible use of generative AI (Fan et al., 2023), e.g., to protect copyright or to curb misinformation. In particular, the traceability of AI-generated images has become increasingly urgent, as many AI products, such as DALL-E (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022), can create highly photorealistic and artistic images that are hard to distinguish from natural photos or human drawings. Unsurprisingly, some AI-generated images have caused false beliefs on social media (Wendling, ). Thus, major tech companies such as Google, OpenAI (Bartz & Hu, ) have recently opted to incorporate watermarks into their image generation products to improve traceability and promote responsible use.
@@
Watermarking methods can be generally divided into two categories: (i) non-blind methods (Cox et al., 1997; Hsieh et al., 2001; Pereira & Pun, 1999) and (ii) blind methods (Bi et al., 2007), which are divided by whether access to original clean images is required to correctly decode the watermarked images (Zhao et al., 2024). In what follows, we focus only on blind methods (and refer to them as watermarks), as they do not require access to clean images and fit better in large-scale application scenarios, such as tracing AI-generated content. Watermarks are typically embedded in images in terms of an overlaid logo or steganography [^1] Steganography: detectable messages embedded in an image but are invisible to human eyes. (Morkel et al., 2005); see Fig. 1 for an example. The embedded watermarks then are typically detected by human eyes or by algorithmic decoders (Voyatzis & Pitas, 1999; Zhu et al., 2018). In early research, various manually designed watermarks were proposed and applied to copyright protection, e.g., visible watermarks (detectable by human eyes) such as an overlaid layer (Kankanhalli et al., 1999) or a color code signature (such as the example in Appendix A); invisible watermarks (detectable by algorithmic decoders) such as Tirkel et al. (1993); Pereira & Pun (2000); Navas et al. (2008). However, these manually designed watermarks are challenged by robustness concerns, i.e., imperceptible corruption, digital editing, or deliberate attacks to the watermarked image can make them undetectable (Mishra, 2022; Zhao et al., 2023); see also Fig. 2. To address these challenges, recent work has shifted to training-based watermarking systems using deep neural networks (DNNs). By incorporating ideas from data augmentation (Mumuni & Mumuni, 2022) and adversarial training (Goodfellow et al., 2014b), the robustness of training-based watermarks against common digital corruptions consistently beats that of manually designed ones (Zhu et al., 2018; Zhang et al., 2019b; Tancik et al., 2020; Jia et al., 2021).
@@
Robustness evaluation of watermark systems often operates under either the white-box or the black-box evasion (i.e., attack) setting (Zhao et al., 2024; An et al., 2024). In white-box evasion, the decoder and the groundtruth watermark are accessible to a dedicated evader (Jiang et al., 2023), which easily leads to a high threat. In contrast, the threat can be greatly reduced by maintaining secrecy of the system (An et al., 2024), i.e., under the black-box model, where the decoder and the groundtruth watermark remain unknown to the evader. Nowadays, it is common for companies not to open-source their generative products, e.g., DALLE-3 from OpenAI, Midjourney, Imagen from Google. Thus, being robust against potential black-box evasions is of higher priority in practical scenarios.
@@ 2 Background and related work
where $BA$ denotes the bitwise accuracy and $\gamma$ is a preset task-dependent threshold (Jiang et al., 2023; Fernandez et al., 2023; Yu et al., 2021). Due to the similarity between Eq. 3 and Eq. 1, most existing work considers image watermarking as a special application of steganography, e.g., Zhu et al. (2018); Tancik et al. (2020); An et al. (2024); Zhao et al. (2024). As a result, the learning formulation in Eq. 2 is also widely adopted in works that only focus on watermarking systems, e.g., Zhang et al. (2019b); Fernandez et al. (2023).
The above watermark methods are post-processing in nature, as the watermark is embedded on any given image $I$ that is already generated. There is an emerging line of in-processing watermark methods that directly modify the image generation process (Zhao et al., 2024; An et al., 2024), including TreeRing watermark (Wen et al., 2023), stable signature (Fernandez et al., 2023), Gaussian shading watermark (Yang et al., 2024), and pseudorandom error-correcting code watermark Gunn et al. (2024). In these in-processing methods, there is no notion of “clean images” $I$ and every generated watermark image exhibits a semantic shift compared to the image generated without the in-process watermarking; see Fig. 4 for an example generated by the TreeRing watermark. In this paper, we focus on post-processing watermark methods due to their flexibility, as they are agnostic to the image generation process.
@@
Robustness of an image watermarking system refers to the extent of the watermark to remain detectable by the decoder $D$ when the watermarked image is manipulated (also called “evaded” if such manipulation is a deliberate attack). Thus, robustness is typically associated with the potential of a watermarking system to be applied to copyright protection and misinformation detection. To stress test the robustness of watermark systems, various watermark evasion techniques have been proposed. These techniques are broadly classified into white-box and black-box evasions, depending on whether any component of the watermark system is known to the evader (An et al., 2024). In this paper, we focus on black-box evasions where nothing about the watermark system is known, as in practice, companies tend to keep their watermark system private. Existing black-box evasions can be classified into two groups: Corruption methods try to distort watermarked images so that watermarks become corrupted and undetectable. Classical digital editing (e.g., applying Gaussian noise, Gaussian blur, JPEG compression, etc. (Voyatzis & Pitas, 1999)), the query-based adversarial attack (WevadeBQ) in Jiang et al. (2023) and the surrogate attack in Saberi et al. (2023) belong to this group; Purification methods treat embedded watermark patterns as noise signals and attempt to remove them using denoising and regeneration techniques, such as BM3D (Dabov et al., 2007), diffusion models (Saberi et al., 2023; Zhao et al., 2023) and VAE (Zhao et al., 2023). The rationale behind the purification methods is rooted in Eq. 3, where the original image, if recovered successfully, is always an evasion. In addition, the original image will be the ultimate threat to any watermarking system—the watermark is evaded by an image without any loss of quality.
@@ 3.1 Watermark evasion via DIP-based blind denoising
4:Query the detection API using all $I_{i}$’s; return all $I_{i}$’s that have no watermark detected.
Now, consider an arbitrary watermarked image $I_{\mathbf{w}}\doteq E(I,\mathbf{w})$, where we do not know $E$ or $\mathbf{w}$. Since $I_{\mathbf{w}}\approx I$ as required in Eq. 3 and $I$ is clearly a successful invasion, it is sensible to try to “purify” or “denoise” $I_{\mathbf{w}}$ toward $I$—the intuition behind all purification methods for black-box evasion (Dabov et al., 2007; Saberi et al., 2023; Zhao et al., 2023).
@@
for any given watermarked image $I_{\mathbf{w}}$. Unlike DIP-based blind denoising which requires appropriate early stopping strategies to find optimal denoising, we only need to check the evasion success of all iterates when iteratively solving Eq. 7 by querying the watermark decoder. Our whole algorithm pipeline is summarized in Algorithm 1. While we do not invent DIP-based blind denoising, we are the first to explore it for blind evasion of invisible watermarks. Rishik also performs DIP-based watermark removal, but solves it as image inpainting which requires known watermark patterns and locations.
@@ 3.2 Why including DIP-based evasion as a baseline method?
3.3 Remarks on improving the query efficiency
Algorithm 1 often takes up to $\sim 1000$ iterative steps to solve Eq. 7. Querying the decoder using all intermediate iterates, as described in Algorithm 1, may be costly, and the decoder server may also limit query frequency for individual users to ensure safety and fairness. To reduce the number of queries and thereby improve query efficiency, it makes sense to only send high-quality iterates to the decoder—again, we emphasize both image quality and evasion success in this paper. To this end, although the clean image $I$ is unknown, the watermarked image $I_{\mathbf{w}}$ is very close to $I$ by design (i.e., per Eq. 3). Therefore, we can take $I_{\mathbf{w}}$ as a proxy reference to estimate the image quality of intermediate iterates; see Appendix C for exemplar trajectories of image quality and watermark detectability on different watermarks by different evasion methods.
@@ 4 Qualitative and quantitative evaluation
(1) Datasets: We use images from two large-scale datasets: (i) MS-COCO (Lin et al., 2014) composed of 328K real images and (ii) DiffusionDB (Wang et al., 2022) composed of 14 million high-quality AI-generated images. We randomly sample $2000$ images from each dataset—the typical scale for the robustness evaluation of watermark systems (An et al., 2024), resize them to $512\times 512$, and generate images with different watermarks, respectively; (2) Watermark methods: We focus on $6$ representative and publicly available post-processing watermark methods: DwtDctSVD, rivaGAN, SSL, TrustMark, RoSteALS, and StegaStamp, whose watermark patterns vary in the visibility level and Fourier spectrum; see Figs. 3 and 6 and Table 1 for visual and quantitative comparisons, respectively. We also evaluate on the SOTA in-processing TreeRing watermark, where $2000$ watermarked images are generated using Gustavosta stable diffusion prompts from HuggingFace (Santana, ); (3) Evasion methods: In addition to our DIP-based evasion method described in Algorithm 1,  [^2] We use the default ‘skip’ network in the original DIP repo: https://github.com/DmitryUlyanov/deep-image-prior. we also consider the following classical digital editing methods: (i) brightness, (ii) contrast, (iii) Gaussian noise, (iv) JPEG compression, (v) bm3d denoising, and recent SOTA purification methods: (vi) DiffPure (Saberi et al., 2023), (vii) Diffuser and (viii) VAE regeneration (Jiang et al., 2023) [^3] VAE regeneration using model from Cheng et al. (2020), which gives the best VAE performance as in Jiang et al. (2023)..
@@
Moreover, comparing the results of our DIP-based evasion on different watermarks, we observe that evasion images for DwtDctSVD, rivaGAN, and SSL watermarks can achieve very high quality when $\gamma\geq 0.65$, with reasonable TPR/FPR values. This is well expected, as these watermarks mainly cause high-frequency distortions (see Fig. 6), and DIP-based evasion is good at separating high- and low-frequency components and thereby largely removing the watermark during iteration, as argued in Section 3.2. In contrast, DIP-based evasion shows limited performance on RoSteALS and StegaStamp, as these watermark methods induce substantial mid- and low-frequency distortions, which DIP picks up in early stages so are hard to separate from the clean image content. DIP’s mediocre performance on TrustMark is a bit surprising. One possible explanation is that TrustMark, in fact, induces mainly low-frequency distortions, which is downplayed in our visualization due to the $\log$ scale as shown in Fig. 6.

### v2

@@ Abstract
Image watermarks have been considered a promising technique to help detect AI-generated content, which can be used to protect copyright or prevent fake image abuse. In this work, we present a black-box method for removing invisible image watermarks, without the need of any dataset of watermarked images or any knowledge about the watermark system. Our approach is simple to implement: given a single watermarked image, we regress it by deep image prior (DIP). We show that from the intermediate steps of DIP one can reliably find an evasion image that can remove invisible watermarks while preserving high image quality. Due to its unique working mechanism and practical effectiveness, we advocate including DIP as a baseline invasion method for benchmarking the robustness of watermarking systems. Finally, by showing the limited ability of DIP and other existing black-box methods in evading training-based visible watermarks, we discuss the positive implications on the practical use of training-based visible watermarks to prevent misinformation abuse. Our code is publicly available at: https://github.com/sun-umn/DIP_Watermark_Evasion_TMLR.
@@ 1 Introduction
In this prosperous era of generative AI, the traceability of AI-generated content (e.g., language, images, and videos) to its source has been frequently mentioned as a promising solution to promote the responsible use of generative AI (Fan et al., 2023), e.g., to protect copyright or to curb misinformation. In particular, the traceability of AI-generated images has become increasingly urgent, as many AI products, such as DALL-E (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022), can create highly photorealistic and artistic images that are hard to distinguish from natural photos or human drawings. Unsurprisingly, some AI-generated images have caused false beliefs on social media (Wendling, ). Thus, major tech companies such as Google, OpenAI (Bartz & Hu, 2023) have recently opted to incorporate watermarks into their image generation products to improve traceability and promote responsible use.
@@
Watermarking methods can be generally divided into two categories: (i) non-blind methods (Cox et al., 1997; Hsieh et al., 2001; Pereira & Pun, 1999) and (ii) blind methods (Bi et al., 2007), which are divided by whether access to original clean images is required to correctly decode the watermarked images (Zhao et al., 2024a). In what follows, we focus only on blind methods (and refer to them as watermarks), as they do not require access to clean images and fit better in large-scale application scenarios, such as tracing AI-generated content. Watermarks are typically embedded in images in terms of an overlaid logo or steganography [^1] Steganography: detectable messages embedded in an image but are invisible to human eyes. (Morkel et al., 2005); see Fig. 1 for an example. The embedded watermarks then are typically detected by human eyes or by algorithmic decoders (Voyatzis & Pitas, 1999; Zhu et al., 2018). In early research, various manually designed watermarks were proposed and applied to copyright protection, e.g., visible watermarks (detectable by human eyes) such as an overlaid layer (Kankanhalli et al., 1999) or a color code signature (such as the example in Appendix A); invisible watermarks (detectable by algorithmic decoders) such as Tirkel et al. (1993); Pereira & Pun (2000); Navas et al. (2008). However, these manually designed watermarks are challenged by robustness concerns, i.e., imperceptible corruption, digital editing, or deliberate attacks to the watermarked image can make them undetectable (Mishra, 2022; Zhao et al., 2024b); see also Fig. 2. To address these challenges, recent work has shifted to training-based watermarking systems using deep neural networks (DNNs). By incorporating ideas from data augmentation (Mumuni & Mumuni, 2022) and adversarial training (Goodfellow et al., 2014b), the robustness of training-based watermarks against common digital corruptions consistently beats that of manually designed ones (Zhu et al., 2018; Zhang et al., 2019b; Tancik et al., 2020; Jia et al., 2021).
@@
Robustness evaluation of watermark systems often operates under either the white-box or the black-box evasion (i.e., attack) setting (Zhao et al., 2024a; An et al., 2024). In white-box evasion, the decoder and the groundtruth watermark are accessible to a dedicated evader (Jiang et al., 2023), which easily leads to a high threat. In contrast, the threat can be greatly reduced by maintaining secrecy of the system (An et al., 2024), i.e., under the black-box model, where the decoder and the groundtruth watermark remain unknown to the evader. Nowadays, it is common for companies not to open-source their generative products, e.g., DALLE-3 from OpenAI, Midjourney, Imagen from Google. Thus, being robust against potential black-box evasions is of higher priority in practical scenarios.
@@ 2 Background and related work
where $BA$ denotes the bitwise accuracy and $\gamma$ is a preset task-dependent threshold (Jiang et al., 2023; Fernandez et al., 2023; Yu et al., 2021). Due to the similarity between Eq. 3 and Eq. 1, most existing work considers image watermarking as a special application of steganography, e.g., Zhu et al. (2018); Tancik et al. (2020); An et al. (2024); Zhao et al. (2024a). As a result, the learning formulation in Eq. 2 is also widely adopted in works that only focus on watermarking systems, e.g., Zhang et al. (2019b); Fernandez et al. (2023).
@@
The above watermark methods are post-processing in nature, as the watermark is embedded on any given image $I$ that is already generated. There is an emerging line of in-processing watermark methods that directly modify the image generation process (Zhao et al., 2024a; An et al., 2024), including TreeRing watermark (Wen et al., 2023), stable signature (Fernandez et al., 2023), Gaussian shading watermark (Yang et al., 2024), and pseudorandom error-correcting code watermark Gunn et al. (2024). In these in-processing methods, there is no notion of “clean images” $I$ and every generated watermark image exhibits a semantic shift compared to the image generated without the in-process watermarking; see Fig. 4 for an example generated by the TreeRing watermark. In this paper, we focus on post-processing watermark methods due to their flexibility, as they are agnostic to the image generation process.
@@
Robustness of an image watermarking system refers to the extent of the watermark to remain detectable by the decoder $D$ when the watermarked image is manipulated (also called “evaded” if such manipulation is a deliberate attack). Thus, robustness is typically associated with the potential of a watermarking system to be applied to copyright protection and misinformation detection. To stress test the robustness of watermark systems, various watermark evasion techniques have been proposed. These techniques are broadly classified into white-box and black-box evasions, depending on whether any component of the watermark system is known to the evader (An et al., 2024). In this paper, we focus on black-box evasions where nothing about the watermark system is known, as in practice, companies tend to keep their watermark system private. Existing black-box evasions can be classified into two groups: Corruption methods try to distort watermarked images so that watermarks become corrupted and undetectable. Classical digital editing (e.g., applying Gaussian noise, Gaussian blur, JPEG compression, etc. (Voyatzis & Pitas, 1999)), the query-based adversarial attack (WevadeBQ) in Jiang et al. (2023) and the surrogate attack in Saberi et al. (2023) belong to this group; Purification methods treat embedded watermark patterns as noise signals and attempt to remove them using denoising and regeneration techniques, such as BM3D (Dabov et al., 2007), diffusion models (Saberi et al., 2023; Zhao et al., 2024b) and VAE (Zhao et al., 2024b). The rationale behind the purification methods is rooted in Eq. 3, where the original image, if recovered successfully, is always an evasion. In addition, the original image will be the ultimate threat to any watermarking system—the watermark is evaded by an image without any loss of quality.
@@ 3.1 Watermark evasion via DIP-based blind denoising
4:Query the detection API using DIP intermediate steps ($I_{i}$’s); return the $I_{i}$’s that have no watermark detected.
Now, consider an arbitrary watermarked image $I_{\mathbf{w}}\doteq E(I,\mathbf{w})$, where we do not know $E$ or $\mathbf{w}$. Since $I_{\mathbf{w}}\approx I$ as required in Eq. 3 and $I$ is clearly a successful invasion, it is sensible to try to “purify” or “denoise” $I_{\mathbf{w}}$ toward $I$—the intuition behind all purification methods for black-box evasion (Dabov et al., 2007; Saberi et al., 2023; Zhao et al., 2024b).
@@
for any given watermarked image $I_{\mathbf{w}}$. Unlike DIP-based blind denoising which requires appropriate early stopping strategies to find optimal denoising, we only need to check the evasion success of all iterates when iteratively solving Eq. 7 by querying the watermark decoder. Our entire algorithm pipeline is summarized in Algorithm 1. Although we do not invent DIP-based blind denoising, we are the first to explore it for blind evasion of invisible watermarks. While Rishik (2020) also performs DIP-based watermark removal, it aims at a different setup: Rishik (2020) tries to remove visible watermarks by DIP-based image inpainting, which requires the exact watermark location as its inpainting mask; in contrast, here we try to remove invisible watermarks using DIP-based denoising, which requires only the watermarked image and nothing else.
@@ 3.2 Why including DIP-based evasion as a baseline method?
3.3 Remarks on DIP-based watermark evasion
Algorithm 1 presents a template algorithm of our DIP-based watermark evasion. Depending on the downstream application, we may need to query a decoder (e.g., a detection API) as stated in Line 3 of Algorithm 1. For example, when the goal is to evade detection while maintaining high image quality, it is essential to query the decoder multiple times to identify the optimal $I_{i}$. In contrast, for single-shot evasion where image quality is less critical, querying the decoder becomes unnecessary. The consideration applies to all existing black-box evasion methods (e.g., those evaluated in Section 4).
On the other hand, querying the decoder in every intermediate iterate, as described in Algorithm 1, may not be possible, especially since the algorithm often requires up to $\sim 1000$ iterations—for example, the decoder server may impose rate limits on individual user to ensure safety and fairness. It is therefore important to reduce the number of decoder queries and improve query efficiency. One potential approach is to selectively query only high-quality iterates—although the clean image $I$ is unknown, the watermarked image $I_{\mathbf{w}}$ is, by design (see Eq. 3), very close to $I$, and thus can serve as a proxy for estimating the image quality of intermediate iterates; see Appendix D for sample trajectories of image quality and watermark detectability across different watermarks and evasion methods.
Finally, although DIP evasion relies on iterative optimization, it remains computationally efficient in practice. For details on its runtime performance, we refer the reader to Appendix C.
@@ 4 Qualitative and quantitative evaluation
(1) Datasets: We use images from two large-scale datasets: (i) MS-COCO (Lin et al., 2014) composed of 328K real images and (ii) DiffusionDB (Wang et al., 2022) composed of 14 million high-quality AI-generated images. We randomly sample $2000$ images from each dataset—the typical scale for the robustness evaluation of watermark systems (An et al., 2024), resize them to $512\times 512$, and generate images with different watermarks, respectively; (2) Watermark methods: We focus on $6$ representative and publicly available post-processing watermark methods: DwtDctSVD, rivaGAN, SSL, TrustMark, RoSteALS, and StegaStamp, whose watermark patterns vary in the visibility level and Fourier spectrum; see Figs. 3 and 6 and Table 1 for visual and quantitative comparisons, respectively. We also evaluate on the SOTA in-processing TreeRing watermark, where $2000$ watermarked images are generated using Gustavosta stable diffusion prompts from HuggingFace (Santana, ); (3) Evasion methods: In addition to our DIP-based evasion method described in Algorithm 1,  [^2] We use the default ‘skip’ network in the original DIP repo: https://github.com/DmitryUlyanov/deep-image-prior. we also consider the following classical digital editing methods: (i) brightness, (ii) contrast, (iii) Gaussian noise, (iv) JPEG compression, (v) bm3d denoising, and recent SOTA purification methods: (vi) DiffPure (Saberi et al., 2023), (vii) Diffuser (Diffusion attack from (Zhao et al., 2024b)) and (viii) VAE regeneration (VAE attack from (Zhao et al., 2024b))  [^3] VAE regeneration using model from Cheng et al. (2020), which gives the best VAE performance as in Jiang et al. (2023)..
@@
Watermarks (Low) Band 1 Band 2 Band 3 Band 4 Band 5 Band 6 Band 7 Band 8 Band 9 Band 10 (High)
dwtDctSvd 0.025 0.044 0.059 0.074 0.091 0.103 0.109 0.109 0.110 0.371
rivaGAN 0.016 0.033 0.061 0.097 0.132 0.163 0.196 0.214 0.237 0.240
SSL 0.011 0.036 0.070 0.119 0.174 0.222 0.289 0.331 0.377 0.765
Trustmark 0.017 0.075 0.177 0.187 0.189 0.175 0.148 0.113 0.094 0.109
Rosteals 0.085 0.282 0.386 0.315 0.411 0.436 0.475 0.556 0.620 0.307
StegaStamp 0.166 0.350 0.432 0.443 0.492 0.518 0.503 0.481 0.434 1.768
Table 5: Average relative Fourier Band Error (FBE) across 10 radial frequency bands for different watermarking methods. Values in bold indicate watermarks with notably high energy in the mid-frequency bands. SSL and Trustmark are also highlighted for comparison: SSL exhibits higher energy in high-frequency components but is vulnerable to DIP evasion, whereas Trustmark shows stronger mid-frequency energy and is more resistant to DIP evasion.
Moreover, comparing the results of our DIP-based evasion on different watermarks, we observe that evasion images for DwtDctSVD, rivaGAN, and SSL watermarks can achieve very high quality when $\gamma\geq 0.65$, with reasonable TPR/FPR values. This is well expected, as these watermarks mainly cause high-frequency distortions (see Fig. 6), and DIP-based evasion is good at separating high- and low-frequency components and thereby largely removing the watermark during iteration, as argued in Section 3.2. In contrast, DIP-based evasion shows limited performance on TrustMark, RoSteALS, and StegaStamp. This is because these watermark methods induce substantial mid- and low-frequency distortions, which DIP picks up in early stages and so are hard to separate from the clean image content.
To better understand this, we compute the average relative FBE across 10 radial frequency bands using 100 images watermarked by each method. This is done by partitioning the Fourier spectrum of the watermark pattern ($I_{w}-I$) into different radial bands—a quantitative analysis of qualitative results in Fig. 6. As shown in Table 5, Trustmark, RoSteALS, and StegaStamp exhibit significantly higher FBE in the lower mid-frequency bands (Bands 3–5) compared to DwtDctSVD, rivaGAN, and SSL. Notably, comparing SSL (whose FBE has larger magnitudes on Band 6-10) and Trustmark (whose FBE has larger magnitudes on Band 3-5), it is even clearer that leveraging lower frequency bands is the key to counter DIP evasions.